a galavano meter having resistance of 50 ohm gives full scale deflection for a current of 0.05 A the length in the meter of a resistance wire of area of croos section 0.00297 that can be used to convert the galavanometer in to ammeter which can read maximum 5A current is(specific resistance of wire 5*10^-7)

Grade:12th pass

2 Answers

Jhashank Dadhich
5 years ago
For ammeter setup formed using galvanometer
(i - ig) Rs  =  (ig) ×(Rg)
(5 - 0.05)Rs = 0.05×50
So  Rs = 0.50 =( rho) L/A
 
(5*10^-7)(L)/0.00297 = 0.50
So L = 2.97*10^-11
Here
Rs = shunt resistor
Rg = resistance of galvanometer
ig = current in galvanometer 
i = total current given 
 
Jhashank Dadhich
5 years ago
You must provide units of all measurements. 
IT is not a correct question. So please provide units also.